OpenStudy (anonymous):

Which of the following definite integrals represents the arc length of the curve y=sqrt(1+x^2) between x = -1 and x = +1 ?

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}\frac{ \sqrt{1+2x^2} }{ \sqrt{1+x^2} }dx\]

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}\frac{ \sqrt{4+5x^2} }{ \sqrt{4+4x^2} }dx\]